fine the slope, x-intercept, and y-intercept for the linear equation 4x+3y=9
Answer by sofiyac(983)

4x+3y=9
3y=-4x+9
y=-4/3x+3 now that the equation is in slope intercept form (y=mx+b) the slope is m so the slope is -4/3
plug in x=0 and you get y=3 so y intercept is (0,3)
plug in y=0 and you get 4x=9 then divide each side by 4 and x=9/4 so x intercept would be (9/4,0)